# GRIP: Algorithm-Agnostic Machine Unlearning for Mixture-of-Experts via Geometric Router Constraints

### v1

@@ 4 Methodology
In this section, we present the GRIP framework. We first motivate this approach through empirical analysis of MoE routing dynamics. Then we formalize routing preservation using null-space constraints (Section 4.1). Building off this, we develop a more flexible, expert-specific constraint decomposition (Section 4.2) and introduce two distinct methods for enforcing these constraints: a training-time projected gradient descent (Section 4.3) and an efficient post-training analytical correction (Section 4.4).
@@
Figure 1(a) confirms that unconstrained optimization preferentially converges to the latter as router drift accumulates across network depth. Crucially, this failure is structural; minor representation drift in shallow layers ($\ell=1\text{--}10$) alters the input manifold for deeper routers, invalidating their learned decision boundaries and triggering a cascading collapse. This observation dictates our design philosophy: effective MoE unlearning requires decoupling routing stability from parameter plasticity. We cannot rely on soft penalties to prevent this structural drift; we must enforce stability geometrically by restricting updates to the subspace where routing decisions remain invariant.
@@ 4.2 Expert-Specific Constraint Decomposition
This expert-specific decomposition significantly expands the feasible optimization landscape compared to global constraints. By treating non-selected experts as inequality constraints rather than hard equalities (enforced via the projection algorithm in Section!4.3), we recover nearly the full gradient space, allowing for robust parameter updates while maintaining strict routing guarantees
4.3 Training-Time Enforcement via Constrained Optimization
@@
Unlike the global constraint, our expert-specific formulation introduces dynamic, input-dependent inequality bounds (Eqs. 45) that preclude enforcement via static reparameterization. Consequently, we employ Projected Gradient Descent (PGD), satisfying equality constraints through subspace projection and inequality constraints via iterative half-space projection.
@@
For expert $j$ at layer $l$, score preservation is required only for the subset of routers selected by the input. We therefore define a local, expert-specific null-space $\mathcal{N}(X^{(j)}_{\text{eq}})$, derived distinctively from the activating representations $X^{(j)}_{\text{eq}}=[x^{(i)}_{r,l}]_{i\in\mathcal{I}_{j,l}}$. By constructing the projector $P_{j,l}$ via the eigenvectors of $X^{(j)}_{\text{eq}}$ (analogous to Section 4.1), we ensure the gradient update remains orthogonal to the subspace governing the routing scores, thereby preserving equality constraints.
To satisfy inequality constraints, we isolate updates that violate the selection margin and project them onto the feasible half-space $\mathcal{H}_{i,j}=\{\nabla:\nabla^{T}x^{(i)}_{r,l}\leq\tau_{i,j}-\varepsilon\}$. Geometrically, this operation removes the gradient component parallel to $x^{(i)}_{r,l}$ responsible for the violation while minimizing deviation from the original optimization direction. We employ Iterative Random Projections (Randomized Kaczmarz) to approximate the intersection of these constraints. By randomly shuffling constraint enforcement order at each step, we avoid cyclic dependencies between correlated half-spaces and accelerate convergence to the feasible region (see Algorithm 1, Appendix C). This guarantees that non-selected experts remain strictly outside the top-$k$ set.
@@ 4.4 Post-Training Analytical Correction
where ${X^{\prime}}^{\dagger}_{r,\ell}={X^{\prime}}^{T}_{r,\ell}(X^{\prime}_{r,\ell}{X^{\prime}}^{T}_{r,\ell}+\lambda I)^{-1}$ is the regularized pseudo-inverse with $\lambda=10^{-6}$ (numerical stability details in Appendix C.5).
This approach fundamentally alters the efficiency profile of the intervention. By replacing iterative per-step projections ($O(Kd^{3})$) with a one-time correction ($O(Ld^{3})$), PTC reduces the computational overhead by orders of magnitude. Although this necessitates caching the initial representations $X_{r,\ell}$, the storage cost $O(LdN_{r})$ is minimal ($2\%$ of the parameter footprint for a 30B model) and can be managed via asynchronous offloading to avoid VRAM contention (detailed complexity analysis in Appendix C.5).
@@ 5.1 Main Results
Table 2: Zero-Shot Utility Loss
@@
Direct Comparison with SEUF
To provide a fair comparison with the state-of-the-art MoE unlearning framework SEUF (Zhuang et al., 2025), we evaluate on the same model architecture (Qwen1.5-MoE-A2.7B-Chat) using WMDP-Cyber (MUSE and WMDP-Bio evaluation is left out as SEUF was not originally evaluated on these datasets). Our framework substantially outperforms SEUF across all metrics. While achieving comparable forget accuracy (0.25–0.27 vs 0.25), we demonstrate superior retain accuracy (0.610.63 vs 0.53, a 15–19% improvement) and routing stability (0.940.99 vs 0.87, an 8–14% improvement).
@@
Table 4: Expert forcing vulnerability
[…]
5.2 Adversarial Evaluation via Expert Forcing
@@ 5.3 Ablation Studies and Design Choices
To validate our design, we conducted an ablation study on different constraint formulations, with results shown in Table 4. The full null-space constraint, while achieving high routing stability (RS=0.96), impedes unlearning effectiveness (FA=0.45). Our proposed expert-specific formulation achieves the a better balance, delivering superior unlearning (FA=0.24) and routing stability (RS=0.96), validating its design. Post-training correction trades slight unlearning performance for near-perfect routing stability and computational efficiency.
Further, in comparing the threshold values $\epsilon$, we tested values across four orders of magnitude $(\epsilon\in\{10^{-4},10^{-3},10^{-2},10^{-1}\})$. Values orders of magnitude less than $10^{-2}$ proved overly restricted and resulted in empty null spaces in later layers. Detailed analysis is contained in Appendix D.1
5.3.1 Implications for Security and Robustness
Vulnerability to Side-Channel Routing Inference.
Hardening Against Activation Steering.
@@ Appendix A Implementation Details
A.1 Evaluation Protocols
We conduct a comprehensive evaluation across two established benchmarks: WMDP [23] for hazardous knowledge removal in bio/cyber domains and MUSE [31] for fictional knowledge unlearning with Harry Potter contexts. Following established conventions, WMDP training continues until models approached random performance (25% accuracy) on forget sets or could no longer achieve lower accuracy. For MUSE, we follow the original four-metric protocol (C1-C4) and unlearn until performance on C1-C3 matches that of the original dense models, ensuring a fair comparison. Our evaluation framework measures three critical dimensions: forget accuracy (FA, lower is better), retain accuracy (RA, higher is better), and routing stability (RS, Jaccard similarity between pre- and post-unlearning expert selections, higher is better). The routing stability metric is calculated through Equation  (2) and serves as our primary indicator of expert selection shift mitigation, with values near 1.0 indicating no disruption to the model’s routing architecture.
@@ B.3 Computational Complexity
Appendix C Randomized Kaczmarz for Halfspace Projection

### v2

@@ 4 Methodology
Standard gradient descent preferentially converges to the latter, creating an optimization shortcut. The gradient with respect to the router weights $\nabla_{\Theta}\mathcal{L}$ typically exhibits a larger expected norm than gradients for deep expert weights $\nabla_{E}\mathcal{L}$ due to the vanishing gradient effect through the sparse conditional computation graph. Furthermore, the router’s decision boundary is locally linear with respect to $\Theta$, whereas expert outputs are highly non-linear deep functions. Consequently, a small perturbation $\|\Delta\Theta\|\ll\|\Delta E\|$ is sufficient to flip the selection indicator $\mathbb{I}[j\in\mathcal{S}_{l}(x)]$, effectively ‘hiding’ the forget set with minimal energy, corresponding to a steep descent direction in the loss landscape.
Figure 1(a) confirms this theoretical intuition: unconstrained optimization converges to this routing shortcut as router drift accumulates across network depth. Crucially, this failure is structural; minor representation drift in shallow layers ($\ell=1\text{--}10$) alters the input manifold for deeper routers, invalidating their learned decision boundaries and triggering a cascading collapse. This observation dictates our design philosophy: effective MoE unlearning requires decoupling routing stability from parameter plasticity. Existing attempts to mitigate this via soft regularization, such as SEUF (Zhuang et al., 2025), prove insufficient; the steep gradient descent direction of the routing shortcut easily overwhelms soft penalties. Furthermore, as Zhuang et al. also identify, the alternative of strict parameter rigidity (freezing the router) fails to accommodate the necessary internal reconfigurations of the experts, leading to suboptimal unlearning. Thus, we must enforce stability geometrically by restricting updates to the subspace where routing decisions remain invariant while leaving orthogonal directions free for optimization.
To operationalize this, we present the GRIP framework. We first formalize routing preservation using null-space constraints (Section 4.1). Building off this, we develop a more flexible, expert-specific constraint decomposition (Section 4.2) and introduce two distinct methods for enforcing these constraints: a training-time projected gradient descent (Section 4.3) and an efficient post-training analytical correction (Section 4.4).
@@ 4.2 Expert-Specific Constraint Decomposition
This expert-specific decomposition significantly expands the feasible optimization landscape compared to global constraints. By treating non-selected experts as inequality constraints rather than hard equalities (enforced via the projection algorithm in Section4.3), we recover nearly the full gradient space, allowing for robust parameter updates while maintaining strict routing guarantees
4.3 Training-Time Enforcement via Stochastic Projection
@@
Unlike the global equality constraint, which admits a static closed-form solution, our expert-specific formulation introduces dynamic, input-dependent bounds that preclude enforcement via static reparameterization. Consequently, we employ a Projected Gradient Descent (PGD) framework that addresses the two distinct regimes of our constraints separately.
@@
First, for the subset of inputs that selected expert $j$ ($\mathcal{I}_{j,\ell}$), we require strict score preservation. We enforce this by defining a local, expert-specific null-space $\mathcal{N}(X^{(j)}_{\text{eq}})$, derived from the activating representations $X^{(j)}_{\text{eq}}=[x^{(i)}_{r,\ell}]_{i\in\mathcal{I}_{j,\ell}}$. By constructing the projector $P_{j,\ell}$ via the eigenvectors of $X^{(j)}_{\text{eq}}$ (analogous to Section 4.1), we ensure the gradient update remains strictly orthogonal to the subspace governing the existing routing scores.
However, for the non-selected inputs $x^{(i)}_{r,\ell}$, the router parameters $\Theta$ are constrained only to lie within dynamic half-spaces $\mathcal{H}_{i,j}=\{\Theta:\Theta x^{(i)}_{r,\ell}\leq\tau_{i,j}-\varepsilon\}$. This defines the full feasible region $\mathcal{K}$ as the intersection of the equality subspace and these dynamic inequalities:
$\mathcal{K}=\mathcal{N}(X^{(j)}_{\text{eq}})\cap\bigcap_{i\notin\mathcal{I}_{j,\ell}}\mathcal{H}_{i,j}$ (6)
Because the defining hyperplanes of $\mathcal{H}_{i,j}$ are functions of the input batch itself, $\mathcal{K}$ varies with the data distribution, precluding the construction of a single, static projection operator. Furthermore, computing the exact Euclidean projection onto this polytope at each step requires solving a quadratic program with thousands of constraints ($N_{\text{ineq}}$), creating a prohibitive computational bottleneck. However, the projection onto any single constituent half-space $\mathcal{H}_{i,j}$ admits a closed-form solution. For a violated constraint, we seek the modified gradient $\tilde{\nabla}^{\prime}$ that satisfies the constraint while minimizing deviation from the original optimization direction:
$\tilde{\nabla}^{\prime}=\operatorname*{argmin}_{G}\|G-\tilde{\nabla}\|_{F}^{2}\quad\text{s.t.}\quad\langle G,x^{(i)}_{r,\ell}\rangle=\tau_{i,j}-\varepsilon$ (7)
To render enforcement tractable, we approximate the global projection onto $\mathcal{K}$ via a stochastic process using Randomized Kaczmarz (RK). Rather than solving for the global intersection directly, we sample a single active constraint $i$ at each iteration $t$ with probability proportional to its geometric weight. The gradient is then updated using the closed-form solution to Equation  (7):
$\tilde{\nabla}^{(t+1)}=\tilde{\nabla}^{(t)}-\frac{\max(0,\langle\tilde{\nabla}^{(t)},x^{(i)}_{r,\ell}\rangle-\tau_{i,j}+\varepsilon)}{\|x^{(i)}_{r,\ell}\|^{2}_{2}}x^{(i)}_{r,\ell}$ (8)
This stochastic formulation functions as an unbiased estimator of the true projection while avoiding the cost of full polyhedral enforcement. We rely on the linear convergence guarantee established by Strohmer and Vershynin (2007), which bounds the expected error for consistent linear systems:
$\mathbb{E}[\|\tilde{\nabla}^{(t)}-\tilde{\nabla}^{*}\|_{2}^{2}]\leq(1-\kappa^{-2})^{t}\|\tilde{\nabla}^{(0)}-\tilde{\nabla}^{*}\|_{2}^{2}$ (9)
where $\kappa=\|X\|_{F}\|X^{\dagger}\|_{2}$ is the scaled condition number of the constraint matrix. In the context of MoE unlearning, the high-dimensionality of router inputs ($d\approx 4096$) yields a favorable condition number, ensuring that the error term decays exponentially. This allows us to enforce thousands of expert-specific constraints with high precision using only a small, fixed budget of iterations ($k_{max}$), avoiding the prohibitive computational cost of full second-order optimization.
@@ 4.4 Post-Training Analytical Correction
where ${X^{\prime}}^{\dagger}_{r,\ell}={X^{\prime}}^{T}_{r,\ell}(X^{\prime}_{r,\ell}{X^{\prime}}^{T}_{r,\ell}+\lambda I)^{-1}$ is the regularized pseudo-inverse with $\lambda=10^{-6}$ (numerical stability details in Appendix D.5).
The efficacy of this closed-form correction relies on the hypothesis that unlearning updates, while semantically significant, induce bounded perturbations in the representation space. Our empirical results (specifically the high recovery of Routing Stability in Table 1) validate this assumption, confirming that the representation drift $X_{r,\ell}\to X^{\prime}_{r,\ell}$ sufficiently preserves the local topology to admit a linear realignment correction. While the global mapping of the model is highly non-linear, the representation drift $X_{r,\ell}\to X^{\prime}_{r,\ell}$ preserves the local topology of the data manifold. Consequently, the router’s decision boundaries, linear hyperplanes in $\mathbb{R}^{d}$, do not need to be re-learned from scratch. Instead, they simply require geometric realignment to match the shifted input distribution.
This formulation fundamentally alters the efficiency profile by decoupling unlearning iterations ($K$) from projection operations. While training-time enforcement necessitates projections at every step, scaling as $\mathcal{O}(KLd^{3})$, PTC requires only a single analytical correction per layer, reducing complexity to $\mathcal{O}(Ld^{3})$. Since unlearning typically requires $K\geq 2000$ steps, this reduces the dominant overhead by three orders of magnitude, lowering the training penalty from $1.67\times$ to a negligible $1.21\times$. The associated storage cost $\mathcal{O}(LdN_{r})$ is minor ($\approx 0.5$ GB for a 30B model). Furthermore, because the representations are static, we eliminate VRAM contention via an asynchronous pre-fetching strategy that streams $X_{r,\ell}$ from host memory only during the correction step.
We now proceed to an empirical validation of this post-training correction alongside the training-time enforcement (Section 4.3) to assess their efficacy in mitigating expert selection shift and preserving model utility.
@@ 5.1 Main Results
Table 2: Zero-Shot Utility Loss on General Benchmarks
@@
5.2 Direct Comparison with SEUF
To provide a fair comparison with the state-of-the-art MoE unlearning framework SEUF (Zhuang et al., 2025), we evaluate on the same model architecture (Qwen1.5-MoE-A2.7B-Chat) using WMDP-Cyber (MUSE and WMDP-Bio evaluation is left out as SEUF was not originally evaluated on these datasets). As detailed in Appendix F, our framework substantially outperforms SEUF across all metrics. While achieving comparable forget accuracy (0.25–0.27 vs 0.25), we demonstrate superior retain accuracy (0.610.63 vs 0.53, a 15–19% improvement) and routing stability (0.940.99 vs 0.87, an 8–14% improvement).
Table 3: Expert forcing vulnerability on WMDP Dataset
[…]
Baseline 0.26 0.37 4.0×
SEUF 0.25 0.33 2.6×
GRIP 0.24 0.27 1.0×
5.3 Adversarial Evaluation via Expert Forcing
5.3.1 Implications for Security and Robustness
Vulnerability to Side-Channel Routing Inference.
Hardening Against Activation Steering.
@@ 5.4 Ablation Studies and Design Choices
To validate our design, we conducted an ablation study on different constraint formulations (Table 5) and a sensitivity analysis on the null-space threshold $\epsilon$ (Table 7).
Constraint Formulations.
We evaluate our constraint designs in Table 4. The strict full null-space constraint ($\Delta\Theta X_{r}=0$) successfully preserves routing ($\text{RS}=0.96$) but overly restricts parameter plasticity, degrading unlearning effectiveness ($\text{FA}=0.45$). Our Expert-Specific formulation resolves this trade-off, achieving superior erasure ($\text{FA}=0.24$) while maintaining high stability ($\text{RS}=0.96$).
While Post-Training Correction (PTC) yields the highest stability ($\text{RS}=0.99$) and efficiency ($1.21\times$), its unlearning performance lags slightly behind the Expert-Specific approach ($\text{FA}=0.28$). This deficit highlights the critical role of Training-Time Enforcement: by permitting interim router drift, PTC forces expert parameters to optimize against a non-stationary routing distribution, introducing gradient noise. Online enforcement eliminates this drift, providing a stable optimization landscape that allows experts to learn more precise unlearning representations. We therefore recommend the Expert-Specific formulation for maximum efficacy, reserving PTC for resource-constrained deployments.
@@
Table 5: Sensitivity Analysis of Null-Space Threshold $\epsilon$
Threshold ($\epsilon$) RS ($\uparrow$) FA ($\downarrow$) RA ($\uparrow$)
$10^{-4}$ 0.97 0.31 0.59
$10^{-3}$ 0.96 0.28 0.62
$\mathbf{10^{-2}}$ (Default) 0.96 0.24 0.65
$10^{-1}$ 0.89 0.25 0.59
Threshold Sensitivity Analysis.
We further analyzed the sensitivity of the null-space threshold $\epsilon$, testing values across four orders of magnitude ($\epsilon\in\{10^{-4},10^{-3},10^{-2},10^{-1}\}$). We utilize $\epsilon=10^{-2}$ as the default for all experiments. As shown in Table 5, values orders of magnitude less than $10^{-2}$ proved overly restricted, resulting in empty null spaces in later layers and degraded unlearning performance ($\text{FA}=0.31$). Conversely, overly loose thresholds ($10^{-1}$) lead to a drop in routing stability ($0.89$). The chosen value of $10^{-2}$ represents the optimal pareto frontier between stability and plasticity.
These comprehensive results establish our framework as a fundamental advancement for MoE model unlearning, providing consistent improvements across diverse benchmarks, unlearning algorithms, and evaluation metrics, demonstrating the broad applicability and robustness of our approach.
@@ 8 Limitations
Acknowledgements
R. Wei and P. Li are partially supported by the NSF under awards PHY-2117997, IIS-2239565, IIS-2428777, and CCF-2402816; the JPMorgan Chase Faculty Award; and the OpenAI Researcher Access Program Credit, the Nvidia Academic Award, and the Google Academic Award.
@@ Appendix A Implementation Details
A.1 Evaluation Metric Details
WMDP Metrics.
  • Forget Accuracy (FA): The accuracy of the model on multiple-choice questions (MCQA) derived from the hazardous corpora (Bio/Cyber). Lower is better (approaching random guess $\approx 0.25$).
  • Retain Accuracy (RA): The accuracy on MMLU and general knowledge MCQA tasks. Higher is better (approaching original model performance).
MUSE Metrics. We adhere to the four-way categorization defined by [31]:
  • C1 (Verbatim Memorization): Measures the Rouge-L overlap of model completions given prefixes from the target book. Unlearning aims to minimize this (Lower is better).
  • C2 (Knowledge Probing): Accuracy on MCQA probing specific facts about the book’s plot and characters (Lower is better).
  • C3 (Privacy Leakage): Measures the model’s tendency to output specific proper nouns (entity extraction) associated with the copyrighted text (Lower is better).
  • C4 (Local Hallucination): Measures the consistency of the model on unrelated texts. We report the preservation rate (Higher is better).
A.2 Unlearning Baseline Formulations
To ensure fair comparison, we standardized the loss functions for all baseline methods used in Section 5. Let $f_{\theta}(x)$ denote the model output and $\mathcal{L}_{CE}$ denote the cross-entropy loss.
Gradient Ascent (GA). We maximize the loss on the forget set $D_{f}$ while optionally minimizing loss on the retain set $D_{r}$ (simulated via gradient accumulation or replay).
$\mathcal{L}_{GA}=-\mathcal{L}_{CE}(f_{\theta}(x_{f}),y_{f})+\alpha\mathcal{L}_{CE}(f_{\theta}(x_{r}),y_{r})$ (11)
KL Minimization (KL). We minimize the KL divergence between the unlearned model and a fixed reference model (the initial pre-trained model $\theta_{ref}$) on the forget set, encouraging the model to output a uniform or "refusal" distribution.
$\mathcal{L}_{KL}=\text{KL}(f_{\theta_{ref}}(x_{f})||f_{\theta}(x_{f}))+\mathcal{L}_{CE}(f_{\theta}(x_{r}),y_{r})$ (12)
Negative Preference Optimization (NPO). Following [40], we treat the forget data as a negative preference sample. The objective makes the forget likelihood worse than the reference model while remaining bounded:
$\mathcal{L}_{NPO}=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{f}|x_{f})}{\pi_{ref}(y_{f}|x_{f})}-\beta\log\tau\right)$ (13)
where $\beta$ controls the strength of the penalty and $\tau$ is a margin hyperparameter.
Representation Misdirection (RMU). Following [23], we perturb the internal activations $h_{l}(x_{f})$ of the model at layer $l$ to match a random target vector $v_{rand}$, preventing the extraction of semantic knowledge:
[…]
@@ B.3 Computational Complexity
Appendix C Efficient Implementation and Storage
Enforcing geometric constraints on large-scale Mixture-of-Experts (MoE) models presents a unique systems challenge. While the parameter updates are sparse, the constraint matrices are dense, layer-specific, and expert-specific. Storing a dense $d\times d$ projector for every expert across all layers would require prohibitive memory. For a standard 30B model ($d=4096$, $E=128$, $L=60$), naive storage would consume over 2 TB of VRAM.To resolve this, we implement a Tiered Hierarchical Caching Strategy, reducing the active memory footprint to $1\%$ of the model size without stalling the training pipeline.
C.1 Implicit Low-Rank Storage
Instead of instantiating and storing the full dense projector $P_{\mathcal{N}}\in\mathbb{R}^{d\times d}$, we exploit the spectral properties of the router representations. We observe that the "retain" subspace for any single expert is effectively low-rank ($k\ll d$). Consequently, we compute and store only the orthonormal basis of the null space, $U_{\text{null}}\in\mathbb{R}^{d\times k}$, derived from the Singular Value Decomposition (SVD) of the expert’s input covariance (see Eq. 10).The projection operation is then performed implicitly via matrix-vector multiplication in the low-rank factorized form:
$P_{\mathcal{N}}\nabla=(U_{\text{null}}U_{\text{null}}^{T})\nabla=U_{\text{null}}(U_{\text{null}}^{T}\nabla)$ (21)
This reduces the storage complexity per expert from $\mathcal{O}(d^{2})$ to $\mathcal{O}(dk)$. For a typical rank $k\approx 50$ and hidden dimension $d=4096$, this yields a compression ratio of $\approx 80\times$, allowing the structural constraints for thousands of experts to reside efficiently in system RAM or on disk.
C.2 Hierarchical Caching System
To manage the flow of these constraints during the unlearning optimization, we developed a custom PyTorch autograd hook (MemoryEfficientConstraintHook) that implements a three-tier cache replacement policy (LRU):
  • Tier 1: Hot Cache (VRAM). Stores the pre-computed projectors for the currently active experts in GPU memory. Capacity is strictly limited (e.g., $C_{\text{hot}}=8$ experts) to prevent OOM errors during the backward pass.
  • Tier 2: Warm Cache (Host RAM). Stores compressed basis vectors ($U_{\text{null}}$) for frequently accessed experts in CPU memory. This tier serves as a low-latency buffer, avoiding the I/O bottleneck of disk reads.
  • Tier 3: Cold Storage (NVMe/Disk). The complete set of expert constraints is sharded into individual files on disk. These are loaded lazily only when a "cache miss" occurs in the upper tiers.
C.3 Stochastic Fallback Mechanism
To ensure numerical stability without incurring unnecessary computation, our implementation employs an adaptive verification step. During the backward pass, the hook first applies the primary projection. It then checks the residual violation against the inequality constraints (Eq. 5). Only if the violation exceeds a numerical tolerance ($\tau>1e^{-4}$) does the system trigger the Randomized Kaczmarz solver. This "lazy enforcement" ensures that the computationally intensive iterative projection (Algorithm 1) is only invoked for the small subset of gradients that actually threaten the routing stability bounds.
Appendix D Randomized Kaczmarz for Halfspace Projection